OpenStudy (anonymous):

f(x)=2x/x+3. Find f^-1(x)

sam (.sam.):

\[f(x)=\frac{2x}{x+3}\] ley y=2x/(x+3) \[y=\frac{2x}{x+3}\] \[y(x+3)=2x\] \[xy+3y=2x\] \[xy-2x=-3y\] \[x(y-2)=-3y\] \[x=\frac{-3y}{y-2}\] \[f^{-1}(x)=\frac{-3y}{y-2}\]

OpenStudy (anonymous):

thank you)

OpenStudy (anonymous):

but you forgot to write it like f(x)^-1=-3x/x-2

sam (.sam.):

yeah you should correct that